OpenStudy (anonymous):

A biologist is comparing the growth of a population of flies per week to the number of flies a bullfrog will consume per week. She has devised an equation to solve for which day (x) the bullfrog would be able to eat the entire population. The equation is 2x = 3x − 1. Explain to the biologist how she can solve this on a graph using a system of equations.

OpenStudy (chasepit):

U can subtract the x to make it a slope function

OpenStudy (anonymous):

it's 2^x=3x-1! sorry about that

OpenStudy (phi):

They want you to graph both equations y= 2^x and y= 3x+1 and see where they intersect.

OpenStudy (anonymous):

They intersect at 1, right?

OpenStudy (anonymous):

What do I do now? @phi